# Superradiant Bosons Driving Supermassive Black Hole Mergers

### v1

@@ Abstract
Ultralight bosons (ULBs) can form macroscopic superradiant clouds around spinning black holes. We show that for scalar ULB masses $\mu\sim 10^{-22}-10^{-21}$ eV boson cloud dynamical friction drives supermassive black hole (SMBH) final-parsec evolution in $\lesssim 1$ Gyr and suppresses the nanohertz gravitational wave background with turnover. Considering century-monitored OJ287 system, we place novel bounds restricting ULB mass range $\mu\simeq(8.5-22)\times 10^{-22}$ eV independent of any dark matter assumptions and show ULB drag can also efficiently reconcile any future confirmations of the debated orbital decay excess. Forthcoming pulsar timing array data and precise SMBH orbital timings will decisively test this scenario.
preprint: CTPU-PTC-25-18, KEK-QUP-2025-0010, KEK-TH-2719, KEK-Cosmo-0380
Introduction.— Ultralight bosons (ULBs) – hypothetical quantum fields with masses below the eV scale – naturally arises in many extensions of the Standard Model (SM). Prominent examples include the QCD axion Peccei:1977hh ; Weinberg:1977ma ; Wilczek:1977pj ; Arvanitaki:2009fg , “fuzzy” dark matter (DM) candidates with masses $\mu\sim 10^{-22}$ eV Turner:1983he ; Press:1989id ; Sin:1992bg ; Hu:2000ke ; Goodman:2000tg , and majoron or axion-like fields Chikashige:1980ui ; Gelmini:1980re ; Arvanitaki:2009fg . Their small masses can be ensured by approximate global symmetries broken at high energies. With feeble ULB non-gravitational couplings, cosmological and astrophysical probes (see, e.g. Kimball:2023vxk ) can be especially sensitive beyond and complementary to laboratory searches.
@@
Rapidly rotating (Kerr) black holes (BHs) can efficiently amplify bosonic field waves, extracting rotational energy and forming long-lived macroscopic “gravitational atom” (GA) condensates Arvanitaki:2009fg ; Arvanitaki:2010sy ; Brito:2015oca ; Baumann:2019eav . This superradiant amplification is gravitational analogue of observed Dicke superradiance in quantum optics and condensed matter Dicke:1954zz . The resulting boson clouds have rich phenomena, including monochromatic gravitational wave (GW) emission yoshino2014gravitational ; Chan:2022dkt , restricting population of spinning BHs Ng:2020ruv ; Stott:2020gjj ; Davoudiasl:2019nlo as well as inducing level-mixing and ionization transitions in binaries Baumann:2019ztm ; Ding:2020bnl ; Tomaselli:2023ysb . This offers an intriguing window for exploring fundamental physics.
Supermassive BHs (SMBHs) constitute prime timely targets for investigation. Recent James Webb Space Telescope (JWST) observations have uncovered active galactic nuclei at high redshifts of $z\simeq 10$ Matthee:2023utn ; 2023Natur.621…51D , implying rapid SMBH growth, while pulsar-timing arrays (PTAs) report detection of a stochastic GW background in the nanohertz band NANOGrav:2023gor ; EPTA:2023fyk ; Zic:2023gta ; Xu:2023wog . These observations sharpen the long-standing “final-parsec” puzzle, how SMBH binaries efficiently shrink on sub-pc scales in Hubble time Milosavljevic:2002ht .
In this work, we demonstrate that superradiant ULBs can significantly affect SMBH binary dynamics and efficiently drive merger evolution in the final parsec regime. An excellent testbed is offered by an archetypal SMBH binary OJ287 already monitored for over a century. Recent analysis has claimed potential excess in its orbital-decay power 2024ApJ…962L..40C , although its existence remains debated Deb:2025raq . We derive novel limits on ULBs from OJ287 independent of cosmological abundance, and also show superradiant ULBs can readily reconcile discrepancy in its orbital dynamics if confirmed. We discuss implications for GWs and identify distinct novel signatures that can definitively test this scenario. We adopt $c=\hbar=G=1$.
@@
Ultralight boson superradiance.— ULB fields can efficiently seed macroscopic superradiant clouds around spinning BHs when their Compton wavelength $\lambda_{c}=1/\mu$, with $\mu$ being their mass, exceeds Schwarzschild radius $r_{s}=2M_{\rm BH}$ of BH with mass $M_{\rm BH}$. As we show, superradiance clouds can significantly impact SMBH binary merger dynamics.
@@
Let us consider a minimal scenario of massive scalar field $\Phi$ propagating on a Kerr spacetime background. with Lagrangian contributions $\mathcal{L}\supset-\left(\partial\Phi\right)^{2}/2-\mu^{2}\Phi^{2}/2$, neglecting possible self-interactions (see Arvanitaki:2010sy ; Baryakhtar:2020gao ; Omiya:2022gwu ; Chia:2022udn ; Xie:2025npy for their potential effects). The resulting Klein-Gordon equation governing the ULB behavior can be simplified to a Schrödinger-like equation by expanding it with respect to a small “gravitational fine structure constant” $\alpha=M_{\rm BH}\mu<1$. See Supplemental Material for details. In this regime, the Schwarzschild radius of BH is smaller than the Compton wavelength of ULB and the system can be described as “gravitational atom (GA)” resembling eigenfunctions of a hydrogen atom described by quantum numbers $n$, $l$, $m$ as
@@
The superradiant growth rate of an $|nlm\rangle$ state can be computed analytically using the Detweiler approximation Detweiler:1980uk , considering $\alpha\lesssim 0.3$ Brito2015b ; Cannizzaro:2023jle , $\Gamma_{nlm}=2\,\tilde{r}_{+}\,C_{nl}\,g_{lm}(m\,\Omega_{\rm BH}-\omega_{nlm})\,\alpha^{4l+5}$, where $\Omega_{\rm BH}=\chi/2M_{\rm BH}\tilde{r}_{+}$ is the angular velocity of the outer BH horizon at $\tilde{r}_{+}=1+\sqrt{1-\chi^{2}}$ and the dimensionless BH Kerr spin parameter is defined as $\chi=J/M_{\rm BH}^{2}$, where $J$ being the angular momentum. Here, $C_{nl}$ are numerical coefficients and $g_{lm}=g_{lm}(\chi,\alpha,\mu)$ are functions Detweiler:1980uk ; Baumann:2019eav . In a superradiant state with $\Gamma_{nlm}>0$, ULB cloud will exponentially grow until spin $\chi$ becomes insufficient and saturates at $\chi_{\text{sat}}=4m\alpha/(m^{2}+4\alpha^{2})$.
@@
×$\times\left(\frac{1.83\times 10^{10}M_{\odot}}{M_{\rm BH}}\right)^{2}~{},$ (2)
where $A_{nlm}(x,\theta)$ denotes the spatial density distribution, where $x=r/r_{0}$ and $\beta$ is the mass ratio between the boson cloud and BH.
OJ287 dynamics.— An intriguing SMBH system to explore the dynamical effects of superradiance clouds is blazar OJ287. We illustrate our scenario in Fig. LABEL:fig:illustration. OJ287 is an unusually massive and tight BH binary with primary BH mass of $M_{\rm BH}=1.83\times 10^{10}\,M_{\odot}$, secondary BH mass of $M_{\ast}=1.50\times 10^{8}\,M_{\odot}$ and orbital timing measurements collected over $\sim 120$ yr period. Its orbital parameters include semi-major axis $a=0.056\,\mathrm{pc}$, inclination angle $\iota=90^{\circ}$, and eccentricity $e=0.657$ Dey:2018mjg ; 1997ApJ…481L…5V ; Valtonen:2013spa (see also Supplemental Material). The binary is shrinking due to continuous GW emission, expected from general relativity Peters:1964zz
@@
Recently, it has been reported that OJ287 apparent total orbital-decay power $\langle P_{\rm tot}\rangle=(3.66\pm 0.24)\times 10^{41}\,$ seemingly exceeds the quadrupole GW prediction $\langle P_{\rm GW}\rangle=(2.62\pm 0.02)\times 10^{41}\,$ by $\sim 4.3\sigma$ 2024ApJ…962L..40C . Notably, these estimates rely on period $T$ and $\dot{T}$ obtained from a GW-driven model. While this has been challenged in Deb:2025raq , including use of redshifted period, the suggested alternative approach of estimating $\langle P_{\rm tot}\rangle$, $T$ and $\dot{T}$ itself relies on using a common GW formalism. We set novel ULB limits from OJ287 dynamics and show that superradiant ULBs can account for the potential excesses in orbital decay if confirmed.
ULBs of mass $\mathcal{O}(10^{-21})\,\mathrm{eV}$ can form a superradiance cloud around primary SMBH of OJ287 within the current age of the Universe $\tau_{\mathrm{U}}\simeq 13.8\,\mathrm{Gyr}$, considering high initial spin that we assume for simplicity to be $\chi_{\rm ini}=1$ since the spins of most SMBHs are expected to be significant Volonteri:2004cf . This can be seen from the superradiance growth rate of the fastest growing GA state $|211\rangle$

### v2

# Dark Matter–Independent Orbital Decay Bounds on Ultralight Bosons from OJ287

@@ Abstract
Ultralight bosons, predicted in scenarios beyond the Standard Model and viable dark matter (DM) candidates, can form superradiant clouds around spinning black holes influencing their dynamics. Using century-long monitored OJ287 supermassive black hole binary we set first DM–independent, dynamical constraints on their masses $\mu=(8.5-22)\times 10^{-22}$ eV. These dynamical constraints, driven by boson cloud friction, are robust against DM-model uncertainties and offer a novel ultralight boson probe. We show that analogous superradiant dynamics across the cosmic population of supermassive black hole systems could help resolve final-parsec evolution stalling problem and imprint a detectable suppression and break in the gravitational wave background.
I Introduction
@@
Ultralight bosons (ULBs) with masses below the eV-scale naturally arises in many extensions of the Standard Model (SM) and constitute viable dark matter (DM) candidates (see, e.g. Peccei:1977hh ; Weinberg:1977ma ; Wilczek:1977pj ; Arvanitaki:2009fg ; Turner:1983he ; Press:1989id ; Sin:1992bg ; Hu:2000ke ; Goodman:2000tg ; Chikashige:1980ui ; Gelmini:1980re ). With feeble ULB non-gravitational couplings, cosmological and astrophysical probes (see e.g. Kimball:2023vxk ) can be highly sensitive beyond and complementary to laboratory searches. When the wavelength of ULB is much larger than the black hole (BH) horizon, a rapidly rotating BH can efficiently amplify the bosonic field waves extracting rotational energy and forming long-lived macroscopic “gravitational atom” (GA) clouds Arvanitaki:2009fg ; Arvanitaki:2010sy ; Brito:2015oca ; Baumann:2019eav . This superradiant amplification is the gravitational analogue of observed Dicke superradiance in quantum optics and condensed matter Dicke:1954zz . The resulting rich phenomena yoshino2014gravitational ; Chan:2022dkt ; Ng:2020ruv ; Stott:2020gjj ; Davoudiasl:2019nlo ; Baumann:2019ztm ; Ding:2020bnl ; Tomaselli:2023ysb ; Li:2025qyu offer intriguing windows for exploring fundamental physics.
In this work we introduce a novel dynamical probe of ULBs based on their superradiance impact on BH binary orbits, independent of any DM assumptions. Unlike previous ULB superradiance constraints derived from BH spin measurement statistics (e.g. Davoudiasl:2019nlo ; Saha:2022hcd ), our method uses orbital decays as a new dynamical channel to probe the $\mu\sim 10^{-21}$ eV window. We further show that superradiant boson clouds can accelerate supermassive BH binary evolution in the late stage final-parsec regime, where evolution can problematically stall. An ideal system for this test is the OJ287 supermassive BH (SMBH) binary Deb:2025raq , monitored for over a century. Prior analyses of OJ287 crucially relied on new fields constituting DM abundance Alachkar:2022qdt ; 2024ApJ…962L..40C ; Deb:2025raq . Here, we set the first DM–independent dynamical limits on ULBs from SMBH orbital evolution, and identify distinctive gravitational wave background signatures, connecting particle physics, astrophysics, and gravitational wave (GW) cosmology.
II Ultralight boson superradiance
@@
ULB fields can efficiently seed macroscopic superradiant clouds around spinning BHs when their Compton wavelength $\lambda_{c}=1/\mu$, with $\mu$ being their mass, exceeds Schwarzschild radius $r_{s}=2M_{\rm BH}$ of BH with mass $M_{\rm BH}$. As we show, superradiance clouds can significantly impact SMBH binary merger dynamics. We adopt $c=\hbar=G=1$.
@@
Let us consider a minimal scenario of massive scalar field $\Phi$ propagating on a Kerr spacetime background. with Lagrangian contributions $\mathcal{L}\supset-\left(\partial\Phi\right)^{2}/2-\mu^{2}\Phi^{2}/2$, neglecting possible self-interactions (see Arvanitaki:2010sy ; Baryakhtar:2020gao ; Omiya:2022gwu ; Chia:2022udn ; Xie:2025npy ; Lambiase:2025twn for their potential effects). The resulting Klein-Gordon equation governing the ULB behavior can be simplified to a Schrödinger-like equation by expanding it with respect to a small “gravitational fine structure constant” $\alpha=M_{\rm BH}\mu<1$. See App. A for details. In this regime, the Schwarzschild radius of BH is smaller than the Compton wavelength of ULB and the system can be described as “gravitational atom (GA)” resembling eigenfunctions of a hydrogen atom described by quantum numbers $n$, $l$, $m$ as
@@
The superradiant growth rate of an $|nlm\rangle$ state can be computed analytically using the Detweiler approximation Detweiler:1980uk , considering $\alpha\lesssim 0.3$ Brito2015b ; Cannizzaro:2023jle ,
$\Gamma_{nlm}=2\,\tilde{r}_{+}\,C_{nl}\,g_{lm}(m\,\Omega_{\rm BH}-\omega_{nlm})\,\alpha^{4l+5}~,$ (2)
where $\Omega_{\rm BH}=\chi/2M_{\rm BH}\tilde{r}_{+}$ is the angular velocity of the outer BH horizon at $\tilde{r}_{+}=1+\sqrt{1-\chi^{2}}$ and the dimensionless BH Kerr spin parameter is defined as $\chi=J/M_{\rm BH}^{2}$, where $J$ being the angular momentum. Here, $C_{nl}$ are numerical coefficients and $g_{lm}=g_{lm}(\chi,\alpha,\mu)$ are functions Detweiler:1980uk ; Baumann:2019eav . In a superradiant state with $\Gamma_{nlm}>0$, ULB cloud will exponentially grow until spin $\chi$ becomes insufficient and saturates at
$\chi_{\text{sat}}=\frac{4m\alpha}{m^{2}+4\alpha^{2}}~.$ (3)
@@
$\times\left(\frac{1.83\times 10^{10}M_{\odot}}{M_{\rm BH}}\right)^{2}~,$ (4)
where $A_{nlm}(x,\theta)$ denotes the spatial density distribution, where $x=r/r_{0}$ and $\beta$ is the mass ratio between the boson cloud and BH. For the dominant states in this work
$A_{211}(x,\theta)$ $=\frac{1}{64\pi}x^{2}e^{-x}\sin^{2}\theta~,$ (5)
$A_{31-1}(x,\theta)$ $=\frac{1}{6561\pi}(x^{2}-6x)^{2}e^{-2x/3}\sin^{2}\theta~.$ (6)
In the following discussion, we consider the impact of superradiant bosons in OJ287 orbital dynamics as illustration in Fig. 1.
Figure 1: Schematic illustration of secondary BH (gray) of mass $M_{\ast}$ (black) traversing ULB superradiance cloud in $|211\rangle$ state orbiting primary BH of mass $M_{\rm BH}$ and spin $\chi$ in SMBH binary, such as OJ287 system. Primary BH rotation defines the reference axis. The secondary’s position is given by its separation $R_{\ast}$ and orbital phase $\varphi_{\ast}$ measured from this axis. The orbital angular momentum is perpendicular to the primary BH spin.
The density profile of superradiant bosons plays an important role in determining the orbital dynamics of OJ287, we display the corresponding profiles and enclosed cloud mass for parameters relevant to OJ287 ($\alpha=\beta=0.1$) in Fig. 2.
Figure 2: [Top] Density profiles of the $|211\rangle$ and $|31-1\rangle$ GA states for $\alpha=\beta=0.1$ around primary BH in OJ287 system. [Bottom] Ratio of enclosed superradiance cloud mass relative to SMBH $M_{\rm BH}$. The vertical dashed line mark the semi-major axis ($0.056$ pc) of OJ287. The dotted line corresponds to $50\,r_{s}$.
III OJ287 dynamics and novel boson limits
A pristine SMBH system to explore the dynamical effects of superradiance clouds is blazar OJ287. OJ287 is an unusually massive and tight BH binary with primary BH mass of $M_{\rm BH}=1.83\times 10^{10}\,M_{\odot}$, secondary BH mass of $M_{\ast}=1.50\times 10^{8}\,M_{\odot}$ and orbital timing measurements collected over $\sim 120$ yr period. Its orbital parameters include semi-major axis $a=0.056\,\mathrm{pc}$, inclination angle $\iota=90^{\circ}$, and eccentricity $e=0.657$ Dey:2018mjg ; 1997ApJ…481L…5V ; Valtonen:2013spa as illustration in The binary is shrinking due to continuous GW emission, expected from general relativity Peters:1964zz
@@ III OJ287 dynamics and novel boson limits
Recent work claimed an apparent OJ287 total orbital-decay power $\langle P_{\rm tot}\rangle=(3.66\pm 0.24)\times 10^{41}\,$W seemingly exceeds the quadrupole GW prediction $\langle P_{\rm GW}\rangle=(2.62\pm 0.02)\times 10^{41}\,$W by $\sim 4.3\sigma$ 2024ApJ…962L..40C . However, detailed reanalysis finds only a $0.41\sigma$ deviation Deb:2025raq . Remaining conservative, we base our limits on the null-excess result compared to relativity GW predictions and derive the first DM-independent, dynamical constraint on ULBs from OJ287.
ULBs of mass $\mathcal{O}(10^{-21})\,\mathrm{eV}$ can form a superradiance cloud around primary SMBH of OJ287 within the current age of the Universe $\tau_{\mathrm{U}}\simeq 13.8\,\mathrm{Gyr}$, considering high initial spin $\chi_{i}$ as expected of most SMBHs Volonteri:2004cf . This can be seen from the superradiance growth rate of the fastest growing GA state $|211\rangle$